How do you interpret the value of the slope of Isocost line


1
Expert's answer
2021-05-23T19:14:38-0400

An isocost line can be represented mathematically as

"C=wL+rK"

where;

C=production cost

w=price of labor wages

L= labor units

r=price of capital or interest rate

K=units of capital


The value of the slope of the Isocost line is the ratio of price of input factors.

For example in the figure below, the slope of the Isocost line will be expressed as,


Let's consider the figure below, x-axis rep[resents units of labor while y-axis represents units of capital




"Slope\space of\space AB= \frac{price\space of\space labor}{price\space of\space capital}"


"=\frac{w}{r}"


This value remains the same throughout the isocost line